# NUBO: A Transparent Python Package for Bayesian Optimization

### v2

@@ 3.2 Bayesian optimization
Before describing the individual optimization options in detail, we want to illustrate NUBO’s user experience—its easy-to-read and write syntax, flexibility, and modularity—on a simple Bayesian optimization step that can be further divided into four substeps.
@@
First, we define the input space. Here, we want to optimize a six-dimensional objective function that is bounded by the hyper-rectangle $[0,1]^{6}$ specified as bounds, a $2\times 6$ torch.Tensor, where the first row provides the lower bounds and the second row the upper bounds for all six input dimensions. Second, we load the training inputs x_train and the training outputs y_train. This training data can be manually selected or generated using a space-filling design, such as Latin hypercube sampling introduced in Section 3.3. Third, we define and train the Gaussian process implemented in NUBO as discussed in Section 3.1, or set up a custom Gaussian process with GPyTorch. Fourth, we specify an acquisition function that takes the fitted Gaussian process as an argument and chooses an optimization method. In this case, we use the upper confidence bound introduced in Equation 13 and optimize it with the L-BFGS-B algorithm [3] using the single function.
@@
    $>$ import torch    $>$ from nubo.acquisition import UpperConfidenceBound    $>$ from nubo.models import GaussianProcess, fit_gp    $>$ from nubo.optimization import single    $>$ from gpytorch.likelihoods import GaussianLikelihood    $>$    $>$    $>$ bounds = torch.tensor([[0., 0., 0., 0., 0., 0.],    ...                        [1., 1., 1., 1., 1., 1.]])    $>$    $>$ x_train = # load inputs as torch.Tensor    $>$ y_train = # load outputs as torch.Tensor    $>$    $>$ likelihood = GaussianLikelihood()    $>$ gp = GaussianProcess(x_train, y_train, likelihood=likelihood)    $>$ fit_gp(x_train, y_train, gp=gp, likelihood=likelihood)    $>$    $>$ acq = UpperConfidenceBound(gp=gp, beta=4)    $>$ x_new, _ = single(func=acq, method="L-BFGS-B", bounds=bounds)NUBO is very flexible and allows users to swap out individual elements for other options. For example, we can substitute the UpperConfidenceBound acquisition function or the single optimization strategy without changing any other code lines. This makes it easy and fast to tailor Bayesian optimization to specific problems.
@@
The remainder of this section introduces NUBO’s optimization strategies. Figure 3 shows a flowchart that helps users decide on the correct acquisition function and optimizer for their specific problem.
@@
3.2.1 Sequential single-point optimization
@@
In NUBO, we differentiate between two optimization strategies: single-point and multi-point optimization. When using the single-point strategy via the single function, NUBO uses the analytical acquisition functions discussed in Section 2.2 to find the next point to be evaluated by the objective function. The corresponding observation must be gathered before the next iteration of the optimization loop can begin.
@@
The code below shows how the analytical expected improvement (EI) and the analytical upper confidence bound (UCB) can be specified with NUBO. The former takes the best training output to date as the argument y_best, while the latter accepts the trade-off hyper-parameter $\beta$ as the beta argument. For bounded optimization problems with analytical acquisition functions, the optimization method of the single function should be set to method="L-BFGS-B". The arguments num_starts (default 10) and num_samples (default 100) can be set to enable multi-start optimization, where the selected optimization algorithm is run multiple times. Each start is initialized at the best points from a large number of points sampled from a Latin hypercube. This reduces the risk of getting stuck in a local optimum. Section 3.3 introduces Latin hypercube sampling in more detail. The single function only returns the best start and its acquisition value. Sequential single-point optimization can be paired with constrained and mixed optimization, as detailed in this section.
    $>$ from nubo.acquisition import ExpectedImprovement, UpperConfidenceBound    $>$ from nubo.optimization import single    $>$    $>$    $>$ acq = ExpectedImprovement(gp=gp, y_best=torch.max(y_train))    $>$ acq = UpperConfidenceBound(gp=gp, beta=4)    $>$ x_new, _ = single(func=acq, method="L-BFGS-B", bounds=bounds,    ...                   num_starts=5, num_samples=50)
3.2.2 Parallel multi-point optimization
@@
The second optimization strategy is multi-point optimization. This strategy uses the Monte Carlo acquisition functions outlined in Section 2.2 to find multiple points, also called batches, in each iteration of the Bayesian optimization loop. This strategy is particularly beneficial for objective functions that support parallel evaluations, as points can be queried simultaneously, speeding up the optimization process.
@@
NUBO uses the Monte Carlo versions of expected improvement MCExpectedImprovement and upper confidence bound MCUpperConfidenceBound in unison with either the multi_joint or multi_sequential function to compute batches. The two different options for the multi-point optimization strategy are discussed in Section 2.2. In addition to the arguments of the analytical acquisition functions, both Monte Carlo acquisition functions accept the number of Monte Carlo samples to be used to approximate the acquisition function as the samples argument (default 512). For the optimization functions, the number of points to be computed can be passed to the batch_size argument, while the method should be set to "Adam" to enable stochastic optimization via the Adam algorithm [2]. The Adam algorithm can be fine-tuned by setting the learning rate lr (default 0.1) and the number of optimization steps steps (default 100). Parallel multi-point optimization can be paired with asynchronous, constrained, and mixed optimization, as detailed in this section.
@@
    $>$ from nubo.acquisition import MCExpectedImprovement,    ...                              MCUpperConfidenceBound    $>$ from nubo.optimization import multi_joint, multi_sequential    $>$    $>$    $>$ acq = MCExpectedImprovement(gp=gp, y_best=torch.max(y_train),    ...                             samples=256)    $>$ acq = MCUpperConfidenceBound(gp=gp, beta=4, samples=256)    $>$ x_new, _ = multi_joint(func=acq, method="Adam", lr=0.1,    ...                        steps=100, batch_size=4, bounds=bounds)    $>$ x_new, _ = multi_sequential(func=acq, method="Adam", lr=0.1,    ...                             steps=100, batch_size=4, bounds=bounds)To enable the use of deterministic optimizers, such as L-BFGS-B [3] and SLSQP [33], the base samples used to compute the Monte Carlo samples can be fixed by setting fix_base_samples=True (default False).
    $>$ from nubo.acquisition import MCUpperConfidenceBound    $>$ from nubo.optimization import multi_joint, multi_sequential    $>$    $>$    $>$ acq = MCUpperConfidenceBound(gp=gp, beta=4, fix_base_samples=True)    $>$ x_new, _ = multi_joint(func=acq, method="L-BFGS-B",    ...                        batch_size=4, bounds=bounds)    $>$    $>$ acq = MCUpperConfidenceBound(gp=gp, beta=4, fix_base_samples=True)    $>$ x_new, _ = multi_sequential(func=acq, method="L-BFGS-B",    ...                             batch_size=4, bounds=bounds)
3.2.3 Asynchronous optimization
@@
NUBO supports asynchronous optimization, that is, the continuation of the optimization loop while some points are being evaluated by the objective function. In this case, the Monte Carlo acquisition functions MCExpectedImprovement or MCUpperConfidenceBound are used as outlined in Section 2.2. The code snippet below assumes that the two points x_pend are currently being evaluated. To continue the optimization, these points can be fed into the acquisition function by setting x_pending=x_pend and NUBO will take them into account for the subsequent iteration.
@@
    $>$ import torch    $>$ from nubo.acquisition import MCUpperConfidenceBound    $>$ from nubo.optimization import multi_joint, multi_sequential    $>$    $>$    $>$ x_pend = torch.tensor([[0.2, 0.9, 0.8, 0.4, 0.4, 0.1],    ...                        [0.1, 0.3, 0.7, 0.2, 0.1, 0.2]])    $>$ acq = MCUpperConfidenceBound(gp=gp, beta=4, x_pending=x_pend)    $>$ x_new, _ = multi_joint(func=acq, method="Adam",    ...                        batch_size=4, bounds=bounds)    $>$ x_new, _ = multi_sequential(func=acq, method="Adam",    ...                             batch_size=4, bounds=bounds)While Monte Carlo acquisition functions are approximations of the analytical functions, they are mainly used for computing multiple points, where analytical functions are generally intractable. The Monte Carlo approach can also be used for single-point asynchronous optimization by setting batch_size=1.
3.2.4 Constrained optimization
@@
The simple maximization problem in Equation 1 can be extended by including one or more input constraints
@@
In these instances, NUBO allows constrained Bayesian optimization by using the SLSQP algorithm to optimize the acquisition function. Implementing this method requires the additional step of specifying the constraints cons as a dictionary for one constraint or a list of dictionaries for multiple constraints. Each constraint requires two entries. The first is "type" and can either be set to "ineq" for inequality constraints or "eq" for equality constraints. The second is "fun", which takes a function representing the constraint. The optimizer only selects points for which the constraint functions are greater than or equal to zero for inequality constraints and exactly zero for equality constraints. The code snippet below specifies two constraints: The first is an inequality constraint that requires the first two input dimensions to be smaller than or equal to 0.5. The second is an equality constraint that requires dimensions four, five, and six to sum to 1.2442.
@@
    >>> import torch    >>>    >>>    >>> bounds = torch.tensor([[0., 0., 0., 0., 0., 0.],    ...                        [1., 1., 1., 1., 1., 1.]])    >>> cons = [{"type": "ineq", "fun": lambda x: 0.5 - x[0] - x[1]},    ...         {"type": "eq", "fun": lambda x: 1.2442 - x[3] - x[4] - x[5]}]After setting up the input space using the bounds and constraints, the Bayesian optimization loop is similar to before. We need to set the method argument of the optimization function to "SLSQP" and provide the function with the constraints cons.
    $>$ from nubo.acquisition import UpperConfidenceBound    $>$ from nubo.optimization import single    $>$    $>$    $>$ acq = UpperConfidenceBound(gp=gp, beta=4)    $>$ x_new, _ = single(func=acq, method="SLSQP",    ...                   bounds=bounds, constraints=cons)Constrained Bayesian optimization can be used with analytical and Monte Carlo acquisition functions and single-point, multi-point, asynchronous, and mixed optimization, as detailed in this section.
3.2.5 Mixed optimization
Bayesian optimization predominantly focuses on problems with continuous input parameters since the Gaussian process models all input dimensions as continuous variables. However, NUBO supports optimizing mixed input parameter spaces via a workaround. To do this, NUBO first computes all possible combinations of the discrete parameters. Then, it maximizes the acquisition function for all continuous parameters while holding one combination of the discrete parameters fixed. Once the acquisition function is maximized for each possible discrete combination, the best overall solution is returned. This can be very time-consuming for many discrete dimensions or discrete values.
To implement mixed optimization in NUBO, bounds are specified as before, but the discrete dimensions are additionally defined in a dictionary where the keys are the dimensions (starting from zero), and the values are a list of all possible values for the discrete inputs. The code below specifies dimensions one and five as disc.
    >>> import torch    >>>    >>>    >>> bounds = torch.tensor([[0., 0., 0., 0., 0., 0.],    ...                        [1., 1., 1., 1., 1., 1.]])    >>> disc = {0: [0.2, 0.4, 0.6, 0.8],    ...         4: [0.3, 0.6, 0.9]}After setting up the input space specified by the bounds and discrete values, the Bayesian optimization loop is similar to before. We only need to provide the function with the dictionary specifying the discrete dimensions discrete=disc.
@@ 3.2.5 Mixed optimization
    $>$ from nubo.acquisition import UpperConfidenceBound    $>$ from nubo.optimization import single    $>$    $>$    $>$ acq = UpperConfidenceBound(gp=gp, beta=4)    $>$ x_new, _ = single(func=acq, method="L-BFGS-B",    ...                   bounds=bounds, discrete=disc)Mixed Bayesian optimization can be used in unison with analytical and Monte Carlo acquisition functions as well as single-point, multi-point, asynchronous, and constrained optimization, all of which are detailed in this section.
@@ 3.3 Test functions and utilities
NUBO provides a selection of test functions and utilities to make implementing and testing Bayesian optimization algorithms more convenient. The ten test functions were selected from the virtual library of [24] and represent a variety of challenges, such as bowl-shaped, plate-shaped, valley-shaped, uni-modal, and multi-modal functions. The functions can be imported from the nubo.test_functions module and instantiated by providing the number of dimensions (except for the Hartmann function that comes in 3D and 6D versions), the standard deviation of any noise that should be added, and whether the function should be minimized or maximized. These functions are equipped with the following attributes: the number of dimensions dims, the bounds bounds, and the inputs and outputs of the global optimum optimum.
    $>$ from nubo.test_functions import Ackley, Hartmann6D    $>$    $>$    $>$ func = Ackley(dims=5, noise_std=0.1, minimise=False)    $>$ func = Hartmann6D(minimise=False)    $>$ dims = func.dims    $>$ bounds = func.boundsThe gen_inputs function from the nubo.utils module allows us to generate input data that covers the input space efficiently by sampling a larger number of random Latin hypercube designs [29] and returning the design with the largest minimal distance between all points. Figure 4 compares Latin hypercube sampling to random sampling for two input dimensions. While many random points are near each other, points from the Latin hypercube design effectively cover the whole input space by placing only one point in each row and column. The exact position of the point within the selected square is random. The code snippet below generates five points for each input dimension of the Hartmann function initiated above and uses func to compute the corresponding outputs.
    $>$ from nubo.utils import gen_inputs    $>$    $>$    $>$ x_train = gen_inputs(num_points=dims * 5,    ...                      num_dims=dims,    ...                      bounds=bounds)    $>$ y_train = func(x_train)
@@
Finally, we discuss three convenience functions that can be used for data transformation. normalise and unnormalise can be used to scale input data to the unit cube $[0,1]^{d}$ and back to its original domain by providing the bounds of the input space. Furthermore, the outputs can be centered at zero with a standard deviation of one with the standardise function.
    $>$ from nubo.utils import standardise, normalise, unnormalise    $>$    $>$    $>$ x_norm = normalise(x_train, bounds=bounds)    $>$ x_train = unnormalise(x_norm, bounds=bounds)    $>$ y_stand = standardise(y_train)

### v3

@@ 3.2 Bayesian optimisation
Before describing the individual optimisation options in detail, we want to illustrate NUBO’s user experience, that is, its easy-to-read and write syntax, flexibility, and modularity, on a simple Bayesian optimisation step that can be further divided into four substeps.
@@
First, we define the input space. Here, we want to optimise a six-dimensional objective function that is bounded by the hyper-rectangle $0,1]^{6}$ specified as bounds, a $2\times 6$ torch.Tensor, where the first row provides the lower bounds and the second row the upper bounds for all six input dimensions. Second, we load the training inputs x_train and the training outputs y_train. This training data can be selected manually or generated by using a space-filling design, such as Latin hypercube sampling introduced in Section 3.3. Third, we define and train the Gaussian process implemented in NUBO as discussed in Section 3.1, or set up a custom Gaussian process with GPyTorch. Fourth, we specify an acquisition function that takes the fitted Gaussian process as an argument and chooses an optimisation method. In this case, we use the upper confidence bound introduced in Equation 8 and optimise it with the L-BFGS-B algorithm (Zhu et al., 1997) using the single function.
    >>> import torch
    >>> from nubo.acquisition import UpperConfidenceBound
    >>> from nubo.models import GaussianProcess, fit_gp
    >>> from nubo.optimisation import single
    >>> from gpytorch.likelihoods import GaussianLikelihood
    >>> bounds = torch.tensor([[0., 0., 0., 0., 0., 0.],
    ...                        [1., 1., 1., 1., 1., 1.]])
    >>> x_train = # load inputs as torch.Tensor
    >>> y_train = # load outputs as torch.Tensor
    >>> likelihood = GaussianLikelihood()
    >>> gp = GaussianProcess(x_train, y_train, likelihood = likelihood)
    >>> fit_gp(x_train, y_train, gp = gp, likelihood = likelihood)
    >>> acq = UpperConfidenceBound(gp = gp, beta = 4)
    >>> x_new, _ = single(func = acq, method = "L-BFGS-B", bounds = bounds)
@@
NUBO is very flexible and allows the user to swap out individual elements for other options. For example, we can substitute the UpperConfidenceBound acquisition function or the single optimisation strategy without changing any of the other lines of code. This makes it easy and fast to tailor Bayesian optimisation to specific problems.
@@
The remainder of this section introduces NUBO’s optimisation strategies. Figure 3 shows a flowchart that helps users decide on the right acquisition function and optimiser for their specific problem.
@@
3.2.1 Sequential single-point optimisation
@@
In NUBO we differentiate between two optimisation strategies: single-point and multi-point optimisation. When using the single-point strategy via the single function, NUBO uses the analytical acquisition functions discussed in Section 2.2 to find the next point to be evaluated by the objective function. The corresponding observation must be gathered before the next iteration of the optimisation loop can begin.
@@
The code below shows how the analytical expected improvement (EI) and the analytical upper confidence bound (UCB) can be specified with NUBO. The former takes the best training output to date as the argument y_best, while the latter accepts the trade-off hyper-parameter $\beta$ as the beta argument. For bounded optimisation problems with analytical acquisition functions, the optimisation method of the single function should be set to method = "L-BFGS-B" and the arguments num_starts (default 10) and num_samples (default 100) can be set to enable multi-start optimisation, where the selected optimisation algorithm is run multiple times and each start is initialised at the best points from a large number of points sampled from a Latin hypercube. This reduces the risk of getting stuck in a local optimum. Section 3.3 introduces Latin hypercube sampling in more detail. The single function only returns the best start and its acquisition value. Sequential single-point optimisation can be paired with constrained and mixed optimisation, which are all detailed in this section.
    >>> from nubo.acquisition import ExpectedImprovement, UpperConfidenceBound
    >>> from nubo.optimisation import single
    >>> acq = ExpectedImprovement(gp = gp, y_best = torch.max(y_train))
    >>> acq = UpperConfidenceBound(gp = gp, beta = 4)
    >>> x_new, _ = single(func = acq, method = "L-BFGS-B", bounds = bounds,
    ...                   num_starts = 5, num_samples = 50)
3.2.2 Parallel multi-point optimisation
@@
The second optimisation strategy is multi-point optimisation. This strategy uses the Monte Carlo acquisition functions outlined in Section 2.2 to find multiple points, also called batches, in each iteration of the Bayesian optimisation loop. This strategy is particularly beneficial for objective functions that support parallel evaluations as points can be queried simultaneously, speeding up the optimisation process.
@@
NUBO uses the Monte Carlo versions of expected improvement MCExpectedImprovement and upper confidence bound MCUpperConfidenceBound in unison with either the multi_joint or multi_sequential function to compute batches. The two different options for the multi-point optimisation strategy are discussed in Section 2.2. In addition to the arguments of the analytical acquisition functions, both Monte Carlo acquisition functions accept the number of Monte Carlo samples to be used to approximate the acquisition function as the samples argument (default 512). For the optimisation functions, the number of points to be computed can be passed to the batch_size argument, while the method should be set to "Adam" to enable stochastic optimisation via the Adam algorithm (Kingma and Ba, 2014). The Adam algorithm can be fine-tuned by setting the learning rate lr (default 0.1) and the number of optimisation steps steps (default 100). Parallel multi-point optimisation can be paired with asynchronous, constrained, and mixed optimisation, which are all detailed in this section.
    >>> from nubo.acquisition import MCExpectedImprovement,
    ...                              MCUpperConfidenceBound
    >>> from nubo.optimisation import multi_joint, multi_sequential
    >>> acq = MCExpectedImprovement(gp = gp, y_best = torch.max(y_train),
    ...                             samples = 256)
    >>> acq = MCUpperConfidenceBound(gp = gp, beta = 4, samples = 256)
    >>> x_new, _ = multi_joint(func = acq, method = "Adam", lr = 0.1,
    ...                        steps = 100, batch_size = 4, bounds = bounds)
    >>> x_new, _ = multi_sequential(func = acq, method = "Adam", lr = 0.1,
    ...                        steps = 100, batch_size = 4, bounds = bounds)
@@
To enable the use of deterministic optimisers, such as L-BFGS-B (Zhu et al., 1997) and SLSQP (Kraft, 1994), the base samples used to compute the Monte Carlo samples can be fixed by setting fix_base_samples = True (default False).
    >>> from nubo.acquisition import MCUpperConfidenceBound
    >>> from nubo.optimisation import multi_joint, multi_sequential
    >>> acq = MCUpperConfidenceBound(gp=gp, beta=4, fix_base_samples=True)
    >>> x_new, _ = multi_joint(func=acq, method="L-BFGS-B",
    ...                        batch_size=4, bounds=bounds)
    >>> acq = MCUpperConfidenceBound(gp=gp, beta=4, fix_base_samples=True)
    >>> x_new, _ = multi_sequential(func=acq, method="L-BFGS-B",
    ...                        batch_size=4, bounds=bounds)
3.2.3 Asynchronous optimisation
@@
NUBO supports asynchronous optimisation, that is, the continuation of the optimisation loop while some points are being evaluated by the objective function. In this case, the Monte Carlo acquisition functions MCExpectedImprovement or MCUpperConfidenceBound are used as outlined in Section 2.2. The code snippet below assumes that the two points x_pend are currently in the evaluation process. To continue the optimisation, these points can be fed into the acquisition function by setting x_pending = x_pend and NUBO will take them into account for the subsequent iteration.
    >>> import torch
    >>> from nubo.acquisition import MCUpperConfidenceBound
    >>> from nubo.optimisation import multi_joint, multi_sequential
    >>> x_pend = torch.tensor([[0.2, 0.9, 0.8, 0.4, 0.4, 0.1],
    ...                        [0.1, 0.3, 0.7, 0.2, 0.1, 0.2]])
    >>> acq = MCUpperConfidenceBound(gp = gp, beta = 4, x_pending = x_pend)
    >>> x_new, _ = multi_joint(func = acq, method = "Adam",
    ...                        batch_size = 4, bounds = bounds)
    >>> x_new, _ = multi_sequential(func = acq, method = "Adam",
    ...                        batch_size = 4, bounds = bounds)
@@
While Monte Carlo acquisition functions are approximations of the analytical functions, they are mainly used for computing multiple points, where analytical functions are generally intractable. The Monte Carlo approach can also be used for single-point asynchronous optimisation by setting batch_size = 1.
3.2.4 Constrained optimisation
@@
The simple maximisation problem in Equation 1 can be extended by including one or more input constraints
@@
In these instances, NUBO allows constrained Bayesian optimisation by using the SLSQP algorithm to optimise the acquisition function. Implementing this method requires the additional step of specifying the constraints cons as a dictionary for one constraint or a list of dictionaries for multiple constraints. Each constraint requires two entries. The first is "type" and can either be set to "ineq" for inequality constraints or "eq" for equality constraints. The second is "fun", which takes a function representing the constraint. The optimiser only selects points for which the constraint functions are greater than or equal to zero for inequality constraints and exactly zero for equality constraints. The code snippet below specifies two constraints: The first is an inequality constraint that requires the first two input dimensions to be smaller than or equal to 0.5. The second is an equality constraint that requires dimensions four, five, and six to sum to 1.2442.
    >>> import torch
    >>> bounds = torch.tensor([[0., 0., 0., 0., 0., 0.],
    ...                        [1., 1., 1., 1., 1., 1.]])
    >>> cons = [{"type": "ineq", "fun": lambda x: 0.5 - x[0] - x[1]},
    ...         {"type": "eq", "fun": lambda x: 1.2442 - x[3] - x[4] - x[5]}]
@@
After setting up the input space using the bounds and constraints, the Bayesian optimisation loop is similar to before. We need to set the method argument of the optimisation function to "SLSQP" and provide the function with the constraints cons.
    >>> from nubo.acquisition import UpperConfidenceBound
    >>> from nubo.optimisation import single
    >>> acq = UpperConfidenceBound(gp = gp, beta = 4)
    >>> x_new, _ = single(func = acq, method = "SLSQP",
    ...                   bounds = bounds, constraints = cons)
@@
Constrained Bayesian optimisation can be used with analytical and Monte Carlo acquisition functions as well as single-point, multi-point, asynchronous, and mixed optimisation, all of which are detailed in this section.
3.2.5 Mixed optimisation
Bayesian optimisation is predominantly focused on problems with continuous input parameters since the Gaussian process models all input dimensions as continuous variables. However, NUBO supports the optimisation of mixed input parameter spaces via a workaround. To do this, NUBO first computes all possible combinations of the discrete parameters. Then, it maximises the acquisition function for all continuous parameters while holding one combination of the discrete parameters fixed. Once the acquisition function is maximised for each of the possible discrete combinations, the best overall solution is returned. Note that this can be very time-consuming for many discrete dimensions or discrete values.
To implement mixed optimisation in NUBO, bounds are specified as before, but the discrete dimensions are additionally defined in a dictionary where the keys are the dimensions (starting from zero) and the values are a list of all possible values for the discrete inputs. The code below specifies dimensions one and five as disc.
    >>> import torch
    >>> bounds = torch.tensor([[0., 0., 0., 0., 0., 0.],
    ...                        [1., 1., 1., 1., 1., 1.]])
    >>> disc = {0: [0.2, 0.4, 0.6, 0.8],
    ...         4: [0.3, 0.6, 0.9]}
After setting up the input space specified by the bounds and discrete values, the Bayesian optimisation loop is similar to before. We only need to provide the function with the dictionary specifying the discrete dimensions discrete=disc.
    >>> from nubo.acquisition import UpperConfidenceBound
    >>> from nubo.optimisation import single
    >>> acq = UpperConfidenceBound(gp = gp, beta = 4)
    >>> x_new, _ = single(func = acq, method = "L-BFGS-B",
    ...                   bounds = bounds, discrete = disc)
Mixed Bayesian optimisation can be used in unison with analytical and Monte Carlo acquisition functions as well as single-point, multi-point, asynchronous, and constrained optimisation, all of which are detailed in this section.
@@ 3.3 Test functions and utilities
NUBO provides a selection of test functions and utilities to make implementing and testing Bayesian optimisation algorithms more convenient. The ten test functions were selected from the virtual library of Surjanovic and Bingham and represent a variety of challenges, such as bowl-shaped, plate-shaped, valley-shaped, uni-modal, and multi-modal functions. The functions can be imported from the nubo.test_functions module and instantiated by providing the number of dimensions (except for the Hartmann function that comes in 3D and 6D versions), the standard deviation of any noise that should be added, and whether the function should be minimised or maximised. These functions are equipped with the following attributes: the number of dimensions dims, the bounds bounds, and the inputs and outputs of the global optimum optimum.
    >>> from nubo.test_functions import Ackley, Hartmann6D
    >>> func = Ackley(dims = 5, noise_std = 0.1, minimise = False)
    >>> func = Hartmann6D(minimise = False)
    >>> dims = func.dims
    >>> bounds = func.bounds
The gen_inputs function from the nubo.utils module allows us to generate input data that covers the input space efficiently by sampling a larger number of random Latin hypercube designs (McKay et al., 1979) and returning the design with the largest minimal distance between all points. Figure 4 compares Latin hypercube sampling to random sampling for two input dimensions. While many random points are in close proximity to each other, points from the Latin hypercube design cover the whole input space effectively by only placing one point in each row and column. The exact position of the point within the selected square is random. The code snippet below generates five points for each input dimension of the Hartmann function initiated above and uses func to compute the corresponding outputs.
    >>> from nubo.utils import gen_inputs
    >>> x_train = gen_inputs(num_points = dims * 5,
    ...                      num_dims = dims,
    ...                      bounds = bounds)
    >>> y_train = func(x_train)
@@
Finally, we discuss three convenience functions that can be used for data transformation. normalise and unnormalise can be used to scale input data to the unit cube $0,1]^{d}$ and back to its original domain by providing the bounds of the input space. Furthermore, the outputs can be centred at zero with a standard deviation of one with the standardise function.
    >>> from nubo.utils import standardise, normalise, unnormalise
    >>> x_norm = normalise(x_train, bounds = bounds)
    >>> x_train = unnormalise(x_norm, bounds = bounds)
    >>> y_stand = standardise(y_train)